Paloma Co. has four em employee, a as four employees. FICA Social Security tases are npan dicare ta Payroll companys four employees. claims The company is for the week ended August 25. Payroll records shonw the tollowing mation tor the Gross Pay ugh Aug. 18 Da Kiesha Chee 117,600 6,900 1,250 450 400 30 In addition to gross pay, the company must emploeach employee pays the remaining one-third. The company also contributes an extra 8% of each Required Co pay two-thirds of the $60 per employee weekly health insur- gross pay (at no cost to employees) to a pension fund. e the following for the week ended August 25 (round amounts to the nearest cent): employees FICA withholdings for Social Security. put 2. Each employees FICA withholdings for Medicare. 3. Employers FICA taxes for Social Security 4. Employers FICA taxes for Medicare. 5. Employers FUTA taxes. 6. Employers SUTA taxes. 7. Each employees net (take-home) pay 8. Employers total payroll-related expense for each employee.
Answer #1

Hi

Let me know in case any issue and query:

Answers are highlighted in yellow: 6.20% 1.45%) 6.20%) 1.45% 0.60% 2.15% 6) Solution: (60*1/3 1) 2 4) Income tax Healthcare Employee Employee Employer Employer Employer Employer Employee SUTANet Pay 1,598.80 666.15 356.58 319.40 Withholding Insurance Soc. Sec. Medicare Soc. Sec. Medicare Name Dali Trey Kiesha Chee Total Gross $2,000.00 $ 284.00 $ 20.00 $ 900.00 $ 145.00$ 20.00 $ 450.00 $ 39.00 $ 20.00 $ 400.00 $ 30.00 $ 20.00 $3,750.00 FUTA 68.20 55.80 27.90 24.80 29.00 68.20$ 29.00 13.05 55.80 13.05 6.53 27.90 6.53 0.60 5.80 24.80 $ 5.80 2.40 2.15 8.60 $ 176.70 $ 54.38 $3.00 $ 10.75 Working: For other than below with threshold limits, gross pay has been taken as base for calcluation: Dali Social Security (118500-117400)*6.2% = (7000-6900)*0.6% (7000-6900) * 2.1596 = $ 68.20 $ 0.60 $ 2.15 FUTA for Kiesha SUTA for Kiesha = Calculation of Net Pay JA-B-C-D-E Question: 8: Healthcare Employer Employer Employer Employer Pension Employers Total FUTA Insurance Soc. Sec. Medicare Fund Payroll Expense Name Dali Trey Kiesha Chee Total Gross $ 2,000.00 $ 40.00 $ 68.20 $ 29.00 $ $900.00 $ 450.00 40.00 $ 27.90 6.53$0.60$ 2.15 $ 36.00$ 563.18 $ 400.00 $ 3,750.00 $ 160.00 $ 176.70 54.38 $3.00 $ 10.75 $ 300.00 $4,454.83 SUTA $ 160.00 2,297.20 $ 72.00 $1,080.85 40.00 55.80 13.05 40.00 S 24.80 5.80 60 5 32.00 $ 513.60